Find the range, variance, and standard deviation unless the question asks for something different.



Passenger Vehicle Deaths. The number of people killed in each state from passenger vehicle


crashes for a specific year is shown


778 309 1110 324 705


1067 826 76 205 152


218 492 65 186 712


193 262 452 875 82


730 1185 2707 1279 390


305 123 948 343 602


69 451 951 104 985


155 450 2080 565 875


414 981 2786 82 793


214 130 396 620 797

1
Expert's answer
2022-06-06T16:54:40-0400

Sorted Data Set: 65, 69, 76, 82, 82, 104, 123, 130, 152, 155, 186, 193, 205, 214, 218, 262, 305, 309, 324, 343, 390, 396, 414, 450, 451, 452, 492, 565, 602, 620, 705, 712, 730, 778, 793, 797, 826, 875, 875, 948, 951, 981, 985, 1067, 1110, 1185, 1279, 2080, 2707, 2786.



"Range=2786-65=2721"


"mean=\bar{x}=\dfrac{1}{50}\displaystyle\sum_{i=1}^{50}x_i=631.98"


"Var(X)=\sigma^2=\dfrac{1}{50}\sum_{i=1}^{50}(x_i-\bar{x})^2""=348319.0196"


"\sigma=\sqrt{\sigma^2}=\sqrt{348319.0196}\approx590.18558"